# Antidirected trees in directed graphs

### v1

@@ Abstract.
The Komlós-Sárközy-Szemerédi (KSS) theorem establishes that a certain bound on the minimum degree of a graph guarantees it contains all bounded degree trees of the same order. Recently several authors put forward variants of this result, where the tree is of smaller order than the host graph, and the host graph also obeys a maximum degree condition. Also, Kathapurkar and Montgomery extended the KSS theorem to digraphs. We bring these two directions together by establishing minimum and maximum degree bounds for digraphs that ensure the containment of oriented trees of smaller order. Our result is restricted to balanced antidirected trees of bounded degree.
More precisely, we show that for every $\gamma>0$, $c\in\mathbb{R}$, $\ell\geq 2$, sufficiently large $n$ and all ${k\geq\gamma n}$, the following holds for every $n$-vertex digraph $D$ and every balanced antidirected tree $T$ with $k$ arcs whose total maximum degree is bounded by $(\log n)^{c}$. If $D$ has a vertex of outdegree at least ${(1+\gamma)(\ell-1)k}$, a vertex of indegree at least $(1+\gamma)(\ell-1)k$ and minimum semidegree $\delta^{0}(D)\geq\leavevmode\nobreak\ (\frac{\ell}{2\ell-1}+\gamma)k$, then $D$ contains $T$.
@@ 1. Introduction
In view of Theorem 1.3 (a) and in analogy to Conjecture 1.1, one might suspect that every graph of minimum degree exceeding $k/2$ and maximum degree at least $2k$ contains every $k$-edge tree. This was conjectured [^1] A version with $\geq k/2$ replacing $>k/2$ had been conjectured earlier [2], but was disproved by Hyde and Reed [4]. in [11]. Even if it is false, replacing $2k$ with another function of $k$ will make it correct:
@@
Our main result is a version of Theorem 1.3 for digraphs and balanced antidirected trees, i.e. oriented trees without 2-directed paths and with the same number of vertices in each partition class.
@@
Note that for $\ell=2$, the conditions of Theorem 1.5 correspond to the ones in Theorem 1.3 (b), while for $\ell=3$, the minimum degree bound of about $k/2$ from Theorem 1.3 (a) is replaced with a minimum semidegree bound of about $53k/$ 5 in Theorem 1.5. Letting $\ell$ tend to infinity and $\gamma$ to 00, the minimum semidegree bound of Theorem 1.5 tends to $k/2$, and the bound on $\Delta(D)$ goes to infinity. We do not know whether the bound on $\delta^{0}(D)$ in Theorem 1.5 is essentially best possible for $\ell\geq 3$, but for $\ell=2$ it is, by the examples given above (since any graph can be viewed as a digraph). Therefore, our result asymptotically solves Problem 6.13 from [12] for balanced antidirected trees of bounded degree in digraphs.
@@
A quick idea of our proof method is given in Section 1.1. Although our proof strategy only seems to work for balanced antidirected trees of bounded degree, it may be that a variant of Theorem 1.5 holds for other oriented trees. If we forbid all oriented $4$-cycles in the host $D$, this is indeed true, even with better bounds on the degrees and for all $k$ and $n$. Trujillo-Negrete and the last author showed in [13] that every digraph $D$ having no oriented 4-cycles with $\delta^{0}(D)\geq k/2$ and $\Delta(D)\geq k$ contains all oriented trees with $k$ arcs.
We close the introduction by remarking that if we only consider hosts that are oriented graphs instead of considering arbitrary digraphs, it is possible to drop the condition on $\Delta(D)$ entirely from Theorem 1.5. Namely, Zárate-Guerén and the last author showed in [14] that for all $\gamma,c$, sufficiently large $n$ and $k\geq\gamma n$, every oriented graph $D$ on $n$ vertices with $\delta^{0}(D)>(1+\gamma)k/2$ contains every balanced antidirected tree $T$ with $k$ arcs and with $\Delta(T)\leq(\log(n))^{c}$. We reprove this result as Corollary 4.3 in this paper.
For more context, history and open directions, we refer the interested reader to the survey [12] and the recent [11].
1.1. Organisation of the paper and overview of our proof.
@@
In Section 3 we define anticonnected digraphs and components. The notion of anticonnectivity captures the meaning of vertices being connected by an antidirected path, in the same spirit as connectivity in undirected graphs represents being connected by an (undirected) path and strong connectivity in digraphs represents being connected by a directed path. The difference to those settings is that anticonnected components have two types of vertices (just like antidirected paths), and distinct anticonnected components may intersect in certain ways. We show that, under certain degree conditions, any large anticonnected digraph contains a large matching (see Lemma 3.3), we call such a matching an anticonnected matching.
Then, in Section 4, we show how to embed a tree or forest into a digraph $D$, if, after applying diregularity to $D$, the reduced digraph of $D$ has large anticonnected component (see Lemma 4.2). For this, we first find a large antimatching in the anticonnected component and then embed most of the tree in the pairs corresponding to arcs of the antimatching, using only small bits of the tree (“links") to connect parts embedded in distinct arcs of the antimatching.
@@ 4. Embedding Trees in a Large Anticonnected Component
Assume we are about to embed a seed $s$. If $s\neq r(T)$ then by (3), the parent of $s$ is already embedded in some vertex $w$ of $W$ that is typical to the $S$-slice $U^{S}$ of some $U\in\textsc{In}(C)\cap N^{+}(W)$ (resp. $\textsc{Out}(C)\cap N^{-}(W)$), and we embed $s$ in any neighbour of $w$ fulfilling (3). If $s=r(T)$ is a source (sink) of $T$ we choose any $U\in\textsc{In}(C)$ ($U\in\textsc{Out}(C)$) and embed $s$ in any vertex of $U^{S}$ fulfilling (3). Such an embedding is possible by Fact 2.5, and since $|S\cup\bigcup\mathcal{L}|<\varepsilon s$ by (2).
@@
Applying Lemma 4.2 with $\frac{\gamma}{100}$ and $c$ gives $\varepsilon$ and $n^{\prime}_{0}$, and applying Lemma 2.6 with $\varepsilon$ and $m_{0}=\frac{1}{\varepsilon}$ gives $n^{\prime\prime}_{0}$ and $M_{0}$. Set $n_{0}:=\max\{n^{\prime}_{0},n^{\prime\prime}_{0}\}$. By Lemma 4.2 $(a)$, we only need to find an anticonnected component $C$ of the reduced oriented graph $R$ of $D$ such that $|C|\geq(1+\gamma)k_{R}$ (where $k_{R}$ is defined as in Lemma 4.2). Consider any anticonnected component $C$. By Fact 2.7 and by Lemma 3.2 $(a)$, $C$ has the desired size unless there is a cluster $U\in\textsc{In}(C)\cap\textsc{Out}(C)$. So assume this is the case. Note that $N^{+}(U)\subseteq\textsc{In}(C)$ and $N^{-}(U)\subseteq\textsc{Out}(C)$, and since $D$ is an oriented graph, $N^{+}(U)\cap N^{-}(U)=\emptyset$. So $C$ has the desired size. ∎
@@
Choose any $u\in V(D)$ with $\deg^{+}(u)\geq\Delta(D)$. Let $R$ be an $(\varepsilon,100\sqrt{\varepsilon})$-reduced directed graph of $D-u$, with clusters of size $s$. Set $k_{R}:=\frac{k}{n}\cdot|R|$. By Fact 2.7 with $a:=(\frac{\ell-1}{2\ell-1}+\gamma)\frac{k}{n}$, we have

### v2

@@ Abstract.
By the well-known Komlós-Sárközy-Szemerédi (KSS) theorem, any $n$-vertex graph of minimum degree $n/2+o(n)$ contains all bounded degree trees of order $n$. Several variants of this result have been put forward recently, where the tree is allowed to be of smaller order than the host graph. In these results, the host graph also obeys a maximum degree condition, which is necessary. In a different vein, the KSS theorem was generalised to digraphs.
We bring these two directions together by establishing minimum and maximum degree bounds for digraphs that ensure the containment of oriented trees of smaller order. Our result, Theorem 1.5, is restricted to balanced antidirected trees of bounded degree but may hold for other oriented trees as well. The full statement of Theorem 1.5 can be found below, we will limit ourselves here to state three important implications.
First, the result implies that, asymptotically, every large digraph of minimum semidegree above $32k/$ 3 having vertices of out-degree and in-degree above $k$ contains each large balanced antidirected bounded-degree tree with $k$ arcs. These semidegree bounds are essentially best possible, and are identical to bounds in corresponding results for graphs.
Second, Theorem 1.5 implies that asymptotically, every large digraph of minimum semidegree above $53k/$ 5 having vertices of out-degree and in-degree above $2k$ contains each large balanced antidirected bounded-degree tree with $k$ arcs. A bound of $k/2$ instead of $53k/$ 5 on the minimum semidegree would be tight and is known to be sufficient in corresponding results for graphs and for oriented graphs.
Thirdly, our result gives a version of the digraph KSS theorem for smaller trees. We prove that for all $k$ and $\varepsilon$ there is a $C$ such that every large digraph of minimum semidegree above $(1+\varepsilon)k/2$ having vertices of out-degree and in-degree above $Ck$ contains each large balanced antidirected bounded-degree tree with $k$ arcs. An exact version of our result for graphs and trees of any maximum degree has been proved recently by Hyde and Reed.
@@ 1. Introduction
The bounds on $\delta(G)$ and on $\Delta(G)$ in Theorem 1.3 are essentially best possible. In view of Theorem 1.3 (a) and in analogy to Conjecture 1.1, one might suspect that every graph of minimum degree exceeding $k/2$ and maximum degree at least $2k$ contains every $k$-edge tree. This was conjectured [^1] A version with $\geq k/2$ replacing $>k/2$ had been conjectured earlier [2], but was disproved by Hyde and Reed [4]. in [11]. Even if it is false, replacing $2k$ with another function of $k$ will make it correct:
@@
Our main result is a version of Theorem 1.3 for digraphs and balanced antidirected trees (where a balanced antidirected tree is an oriented tree without 2-directed paths and with the same number of vertices in each partition class). It reads as follows.
@@
The bounds from Theorem 1.5 may not seem aesthetically pleasing at first glance, but plugging in the concrete values $\ell=3$ and $\ell=2$ we obtain bounds that are very similar to the ones from Theorem 1.3:
Corollary 1.2.
For every $\gamma>0$ and $c\in\mathbb{R}$ there is $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$ and all $k\geq\gamma n$ every $n$-vertex digraph $D$ fulfilling
  1. (a)
    $\delta^{0}(D)\geq(1+\gamma){3k}/{5}$ and $\Delta(D)\geq 2(1+\gamma)k$; or
  2. (b)
    $\delta^{0}(D)\geq(1+\gamma){2k}/{3}$ and $\Delta(D)\geq(1+\gamma)k$
@@
By the examples given above, and since any graph can be viewed as a digraph, the bound on $\delta^{0}(D)$ in Corollary 1.2(b) is essentially best possible (and thus Corollary 1.2(b) asymptotically solves Problem 6.13 from [12] for balanced antidirected trees of bounded degree in digraphs). For the same reasons, in both parts the bounds on $\Delta(D)$ are essentially best possible, and in part (a) the bound on $\delta^{0}(D)$ needs to be at least about $k/2$
We do not know if Corollary 1.2(a) can be improved in that sense or if in the general digraph setting, a larger minimum semidegree is necessary. For certain classes of digraphs, $k/2$ is sufficient, with better or no bounds on $\Delta(D)$: first, Trujillo-Negrete and the last author showed in [13] that every digraph $D$ with $\delta^{0}(D)\geq k/2$ and $\Delta(D)\geq k$ and without oriented 4-cycles contains all oriented trees with $k$ arcs. Second, Zárate-Guerén and the last author showed in [14] that for all $\gamma,c$, sufficiently large $n$ and $k\geq\gamma n$, every oriented graph $D$ on $n$ vertices with $\delta^{0}(D)>(1+\gamma)k/2$ contains every balanced antidirected tree $T$ with $k$ arcs and with $\Delta(T)\leq(\log(n))^{c}$. The more general approach from the present paper allows us to quickly reprove the latter result as Corollary 5.3. In Section 2.1 we will explain why Corollary 5.3 is much easier to prove than Theorem 1.5.
If $\ell$ tends to infinity and $\gamma$ approaches 00, the minimum semidegree bound of Theorem 1.5 tends to $k/2$, and the bound on $\Delta(D)$ goes to infinity. More precisely, we have the following corollary which follows from Theorem 1.5 by taking $\gamma=\varepsilon/10$ and $\ell=\lceil\frac{10}{\varepsilon^{2}}\rceil$.
Corollary 1.3.
For every $10<\varepsilon<$ 1, $c\in\mathbb{R}$ there is $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$ and all $k\geq\varepsilon n$ every $n$-vertex digraph $D$ with $\delta^{0}(D)\geq(1+\varepsilon){k}/{2}$ and $\Delta(D)\geq k/\varepsilon$ contains every balanced antidirected tree $T$ with $k$ arcs and with $\Delta(T)\leq(\log n)^{c}$.
Corollary 1.3 constitutes an asymptotic version of Theorem 1.4 for large bounded degree trees in digraphs. Just as Theorem 1.4 can be viewed as variant of the KSS theorem for smaller trees, Corollary 1.3 is a variant of the digraph KSS theorem from [5] for smaller trees (for a somewhat restricted class of trees). For more context, history and open directions, we refer the interested reader to the survey [12] and the recent [11].
@@
We close the introduction by remarking that although our proof strategy only seems to work for balanced antidirected trees of bounded degree, it may be that a variant of Theorem 1.5 or of any of its corollaries holds for other oriented trees. A quick idea of our proof method and its limitations is given in Section 2.1.
2. Overview of our proof and organisation of the paper.
2.1. Overview of the proof of Theorem 1.5.
We will rely on the regularity method for digraphs as well as on existing results and methods for cutting up trees and embedding them into regular pairs. This has been done similarly for embedding bounded-degree trees in graphs in several papers, for instance in [2], and it has been done recently for embedding antidirected trees in oriented graphs in [14]. Embedding into edges of the reduced digraph is straightforward for antidirected trees, however, finding the appropriate connections between the edges of the reduced digraph can be difficult. The connections we need are antidirected paths in the reduced digraph that run between the edges we plan to use for the embedding.
In [14], where a similar result to our Corollary 1.2(a) was shown for oriented graphs, the reduced digraph is shown to have a sufficiently large matching whose edges are connected in the above sense (we call such a matching an anticonnected matching). This was possible even with a bound of only about $k/2$ on the minimum semidegree and no requirement of having vertices of large semidegree because the host was an oriented graph which makes the problem much easier. For instance, in an oriented graph $D$, just one vertex of semidegree at least $k/2$ guarantees that $D$ has more than $k$ vertices, while a digraph of minimum semidegree $k/2$ can have as little as $k/2+1$ vertices.
The main challenge in our proof is that we need to find a similar matching structure as the one used in [14], but in a digraph $D$. In order to get a better grip on the possible connections $D$, we explore the anticonnected components of $D$. As hinted at above, the notion of anticonnectivity captures the meaning of vertices being connected by an antidirected path, in the same spirit as connectivity in undirected graphs represents being connected by an (undirected) path and strong connectivity in digraphs represents being connected by a directed path. The difference to both those settings is that anticonnected components have two types of vertices (just like antidirected paths), and distinct anticonnected components may intersect in certain ways. Consider, for example, this digraph: , which consists of two anticonnected components that intersect at a single vertex.
Using the lower bound on the minimum semidegree from Theorem 1.5, we show that the anticonnected components of $D$ each contain a suitable anticonnected matching. Our plan is to embed the tree into these matchings.
However, each of the anticonnected components of $D$ could be too small to host the tree $T$ on its own. This is where the vertex $u$ of large out- or in-degree comes into play. We prove that $u$ has a non-negligible degree into at least $\ell$ anticonnected components of the reduced digraph (more precisely, $u$ will either send or receive some $\varepsilon k$ arcs to/from suitable vertices of these components). We then cut the tree $T$ at a strategically chosen vertex $x$, embed $x$ into $u$, group the trees of $T-x$ into $\ell$ families of certain total sizes, and embed each family into one of the anticonnected components seen by $u$. To be exact, we first find $x$ and then, depending on $x$, decide whether we take $u$ as the vertex of large out-degree or as the vertex of large in-degree of $D$.
Unfortunately, there is still one more obstacle to overcome: the components seen by $u$ could intersect. Thus after embedding part of the tree into the first components, parts of the other components may have become unusable. To address this issue, we have to take care to group the trees of $T-x$ into families of the correct sizes, and we will choose the anticonnected components we use dynamically.
Let us now address some of the limitations of our proof. First, our trees need to be antidirected, as otherwise it would not be possible to use the regular pairs for the embedding. Second, our trees need to be balanced as this allows us to control how we fill the clusters of the anticonnected components. Finally our trees need to have bounded degree because of our strategy of using the connecting paths between matching edges. All these restrictions (antidirected, balanced, bounded degree) come from our proof method and it is not impossible that they could be overcome with different methods.
On the other hand, the requirement of $\Delta(D)\geq k$ is necessary, as otherwise the digraph could be too small to host the tree. In the case of $\delta^{0}(D)$ being bounded asymptotically by $k/2$, vertices of out-/in-degree at least $2k$ become necessary because of examples from [2]. These examples also show that the bounds on $\delta^{0}(D)$ in Corollary 1.2(b) and in Corollary 1.3 are best possible, while those in Corollary 1.2(a) may be improvable.
2.2. Organisation of the paper.
@@
In Section 4 we define anticonnected digraphs and components. We show that, under certain degree conditions, any large anticonnected digraph contains a large matching (see Lemma 4.3), we call such a matching an anticonnected matching.
Then, in Section 5, we show how to embed a tree or forest into a digraph $D$, if, after applying diregularity to $D$, the reduced digraph of $D$ has large anticonnected component (see Lemma 5.2). For this, we first find a large antimatching in the anticonnected component and then embed most of the tree in the pairs corresponding to arcs of the antimatching, using only small bits of the tree to connect parts embedded in distinct arcs of the antimatching.
@@ 5. Embedding Trees in a Large Anticonnected Component
Assume we are about to embed a seed $s$. If $s\neq r(T)$ then by (3), the parent of $s$ is already embedded in some vertex $w$ of $W$ that is typical to the $S$-slice $U^{S}$ of some $U\in\textsc{In}(C)\cap N^{+}(W)$ (resp. $\textsc{Out}(C)\cap N^{-}(W)$), and we embed $s$ in any neighbour of $w$ fulfilling (3). If $s=r(T)$ is a source (sink) of $T$ we choose any $U\in\textsc{In}(C)$ ($U\in\textsc{Out}(C)$) and embed $s$ in any vertex of $U^{S}$ fulfilling (3). Such an embedding is possible by Fact 3.5, and since $|S\cup\bigcup\mathcal{L}|<\varepsilon s$ by (2).
@@
Applying Lemma 5.2 with $\frac{\gamma}{100}$ and $c$ gives $\varepsilon$ and $n^{\prime}_{0}$, and applying Lemma 3.6 with $\varepsilon$ and $m_{0}=\frac{1}{\varepsilon}$ gives $n^{\prime\prime}_{0}$ and $M_{0}$. Set $n_{0}:=\max\{n^{\prime}_{0},n^{\prime\prime}_{0}\}$. By Lemma 5.2 $(a)$, we only need to find an anticonnected component $C$ of the reduced oriented graph $R$ of $D$ such that $|C|\geq(1+\gamma)k_{R}$ (where $k_{R}$ is defined as in Lemma 5.2). Consider any anticonnected component $C$. By Fact 3.7 and by Lemma 4.2 $(a)$, $C$ has the desired size unless there is a cluster $U\in\textsc{In}(C)\cap\textsc{Out}(C)$. So assume this is the case. Note that $N^{+}(U)\subseteq\textsc{In}(C)$ and $N^{-}(U)\subseteq\textsc{Out}(C)$, and since $D$ is an oriented graph, $N^{+}(U)\cap N^{-}(U)=\emptyset$. So $C$ has the desired size. ∎
@@
Choose any $u\in V(D)$ with $\deg^{+}(u)\geq\Delta(D)$. Let $R$ be an $(\varepsilon,100\sqrt{\varepsilon})$-reduced directed graph of $D-u$, with clusters of size $s$. Set $k_{R}:=\frac{k}{n}\cdot|R|$. By Fact 3.7 with $a:=(\frac{\ell-1}{2\ell-1}+\gamma)\frac{k}{n}$, we have